# Origin and emergent features of many-body dynamical localization

### v2

@@ Abstract
Fig. 4(a) shows the GFD $D_{3}$ as a function of $g$ and $K$. It is extracted from a linear fit of the type $\overline{S}_{\beta}=D_{\beta}\mathrm{ln}\mathcal{N}+b_{\beta}$ [46]. Here, $b_{\beta}$ is a subleading correction at finite dimension. Negative $b_{\beta}$ can be related to a nonergodic volume $\Lambda_{\beta}=\mathrm{exp}\{-b_{\beta}\}$ [50]. $\overline{S}_{\beta}$ denotes the averaged $S_{\beta}$ over $500$ eigenstates. One can notice three regimes. We first focus on the small kick strength $K=1$, as shown in Fig. 4(c). When $g$ is small, $D_{3}\gtrsim 0$, thus the Floquet eigenstate is highly localized at regime I. Further increasing $g$ in regime II, $D_{3}$ exhibits a dramatic increase, while the fitted intercept $b_{3}$ drops sharply and then levels off near zero. In this regime, the Floquet eigenstates are extended ($0.4\lesssim D_{3}<1$) but remain nonergodic. As $g$ approaches regime III, $D_{3}$ decreases. This reduction is attributed to the fermionization of the bosons as $g$ approaches infinity, at which point the system behaves like a TG gas [36, 34, 33, 35]. In this case, the eigenstates are nonergodic across all values of $g$ and the system is in the MBDL phase. The non-monotonic change of the GFD aligns with the crossover of the algebraic tail of $W_{\mathcal{I}\mathcal{I}^{\prime}}$, and is also consistent with the recently predicted non-monotonic behavior of the interaction energy in the localized regime [32]. A similar behavior is found for $\beta=1,2$. We also notice that $D_{3}\neq D_{2}\neq D_{1}$, thus the MBDL phase is multifractal. For larger kick strengths, $D_{3}$ in regime II greatly increases towards the ergodic value, as shown in Fig. 4(a). This suggests the occurrence of delocalization at intermediate interaction strengths, where the algebraic hopping amplitude of the NN-dimensional lattice reaches a maximum, as we have anticipated.
@@
Discussions—In summary, we have investigated the microscopic origin of the MBDL phase in the kicked LL model. We find the existence of dynamical localization and its breakdown in different regions of parameter space. Finite interaction strengths introduce an algebraic tail to the localization properties, observed in the mapped model. The exponent and amplitude of this algebraic tail exhibit a nontrivial dependence on the interaction strength, reflecting the characteristic features of the MBDL phase. Further analysis of the many-body eigenstates suggests that MBDL possesses multifractal properties and that finite interactions render the system non-integrable. Our study provides an explanation for the longstanding question of MBDL’s existence at finite interaction strengths and is anticipated to be valid for systems with many particles. In End Matter, we propose an experimental implementation to probe the crossover of the algebraic tail. Our results demonstrate that this phenomenon can be observed in the momentum distribution by quenching cold atoms from the non-interacting to the finite-interaction regime.
@@
Acknowledgment—We thank Prof. Chushun Tian for many helpful suggestions. We also thank Prof. Marcos Rigol for discussions. This work was supported by the National Natural Science Foundation of China (Grants No. 12375021 and No. 12247101), the Zhejiang Provincial Natural Science Foundation of China (Grant No. LD25A050002), the National Key Research and Development Program of China (Grant No. No. 2022YFA1404203) and the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2024-jdzx06). The Innsbruck team acknowledges funding by a Wittgenstein prize grant with the Austrian Science Fund (FWF) project number Z336-N36, by the European Research Council (ERC) with project numbers 789017 and 101201611, by an FFG infrastructure grant with project number FO999896041, and by the FWF’s COE 1 and quantA. Y.G. is supported by the FWF with project number 10.55776/COE1. Z.C. also acknowledges the support from the University of Rochester.
Note—As we were finalizing our paper, we found that H. Olsen et al. [59] had independently posted a preprint on the same topic.

### v3

@@ Abstract
Fig. 4(a) shows the GFD $D_{3}$ as a function of $g$ and $K$. It is extracted from a linear fit of the type $\overline{S}_{\beta}=D_{\beta}\mathrm{ln}\mathcal{N}+b_{\beta}$ [46]. Here, $b_{\beta}$ is a subleading correction at finite dimension. Negative $b_{\beta}$ can be related to a nonergodic volume $\Lambda_{\beta}=\mathrm{exp}\{-b_{\beta}\}$ [54]. $\overline{S}_{\beta}$ denotes the averaged $S_{\beta}$ over $500$ eigenstates. One can notice three regimes. We first focus on the small kick strength $K=1$, as shown in Fig. 4(c). When $g$ is small, $D_{3}\gtrsim 0$, the Floquet eigenstate is thus highly localized at regime I. Further increasing $g$ in regime II, $D_{3}$ exhibits a dramatic increase, while the fitted intercept $b_{3}$ drops sharply and then levels off near zero. In this regime, the Floquet eigenstates are extended ($0.4\lesssim D_{3}<1$) but remain nonergodic. As $g$ approaches regime III, $D_{3}$ decreases. This reduction is attributed to the fermionization of the bosons as $g$ approaches infinity, at which point the system behaves like a TG gas [36, 34, 33, 35]. In this case, the eigenstates are nonergodic across all values of $g$ and the system is in the MBDL phase. The nonmonotonic change of the GFD aligns with the crossover of the algebraic tail of $W_{\mathcal{I}\mathcal{I}^{\prime}}$, and is also consistent with the recently predicted nonmonotonic behavior of the interaction energy in the localized regime [32]. As we have shown in Fig. 3(c), this nonmonotonic change is already present in the mapped model’ eigenstates. This indicates a strong link between the structure of the mapped model and the integrability of the kicked LL model. A similar behavior is found for $\beta=1,2$. We also notice that $D_{3}\neq D_{2}\neq D_{1}$, thus the MBDL phase is multifractal. For larger kick strengths, $D_{3}$ in regime II greatly increases toward the ergodic value, as shown in Fig. 4(a). This suggests the occurrence of delocalization at intermediate interaction strengths, where the algebraic hopping amplitude of the NN-dimensional lattice reaches a maximum, as we have anticipated.
@@
Discussions—In summary, we have investigated the microscopic origin of the MBDL phase in the kicked LL model. We find the existence of dynamical localization and its breakdown in different regions of parameter space. Finite interaction strengths introduce an algebraic tail to the localization properties, observed in the mapped model. The exponent and amplitude of this algebraic tail exhibit a nontrivial dependence on the interaction strength, reflecting the characteristic features of the MBDL phase. Further analysis of the many-body eigenstates suggests that MBDL possesses multifractal properties and that finite interactions render the system nonintegrable. Our study provides an explanation for the long-standing question of MBDL’s existence at finite interaction strengths and is anticipated to be valid for systems with many particles. In End Matter, we propose an experimental implementation to probe the crossover of the algebraic tail. Our results demonstrate that this phenomenon can be observed in the momentum distribution by quenching cold atoms from the noninteracting to the finite-interaction regime.
@@
Acknowledgments—We thank Prof. Chushun Tian for many helpful suggestions. We also thank Prof. Marcos Rigol for discussions. This work was supported by the National Natural Science Foundation of China (Grants No. 12375021 and No. 12247101), the Zhejiang Provincial Natural Science Foundation of China (Grant No. LD25A050002), the National Key Research and Development Program of China (Grant No. 2022YFA1404203), and the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2024-jdzx06). The Innsbruck team acknowledges funding by a Wittgenstein Prize grant with the Austrian Science Fund (FWF) [Grant DOI: 10.55776/Z336] Project No. Z336-N36, by the European Research Council (ERC) with Projects No. 789017 and No. 101201611; by an FFG infrastructure grant with Project No. FO999896041; and by the Austrian Science Fund (FWF)’s [Grant DOI: 10.55776/COE 1] and quantA. Z.C. also acknowledges the support from the University of Rochester.
Data availability—The data that support the findings of this article are openly available [64].
Note added—As we were finalizing our manuscript, we found that H. Olsen et al. [65] had independently posted a preprint on the same topic.
@@
Here, we show the exponent and amplitude of the algebraic tail as a function of $g$ in Fig. S1. Different from using $\gamma$ as unit in the main text, the crossover point of $\lambda$ for different $N$ is slightly different. This also indicates that the effective interaction is affected by density.
@@
Figure S1: The decay exponent $\lambda$ (a) and the amplitude AA (b) as a function of $g$ for different NN.
@@
We can understand this universal behavior of the off-diagonal term on a general basis. At small $K$ limit and using the field operator $\Psi(x)$, the off-diagonal term reads equally
$W_{\mathcal{I}\mathcal{I}^{\prime}}$ $=$ $\langle\mathcal{I}\lvert\tan{\Big[\int\frac{K}{2\hbar_{\mathrm{eff}}}\cos{x}\Psi^{\dagger}(x)\Psi(x)dx\Big]}\rvert\mathcal{I}^{\prime}\rangle\approx\int\frac{K}{2\hbar_{\mathrm{eff}}}\cos{x}\langle\mathcal{I}\lvert\Psi^{\dagger}(x)\Psi(x)\rvert\mathcal{I}^{\prime}\rangle dx=$ (S61)
$=$ $\frac{K}{2\hbar_{\mathrm{eff}}}\int\cos{x}\langle\mathcal{I}\lvert e^{i\hat{P}x}\Psi^{\dagger}(0)\Psi(0)e^{-i\hat{P}x}\rvert\mathcal{I}^{\prime}\rangle dx=\frac{K}{2\hbar_{\mathrm{eff}}}\langle\mathcal{I}\lvert\Psi^{\dagger}(0)\Psi(0)\rvert\mathcal{I}^{\prime}\rangle\int\cos{x}e^{i(Q-Q^{\prime})x}dx,$
where $\hat{P}$ is the total momentum operator, generator of translations, and $Q(Q^{\prime})$ denote the total momentum of state $\mathcal{I}(\mathcal{I}^{\prime})$. One can notice that the integral is expressing momentum conservation in the process, making sure that only states with $|Q-Q^{\prime}|=1$ give a non zero matrix element. The term $\langle\mathcal{I}\lvert\Psi^{\dagger}(0)\Psi(0)\rvert\mathcal{I}^{\prime}\rangle$ is nothing but the form factor of the density operator $\Psi^{\dagger}(0)\Psi(0)$ in the integrable LL model. A well-known exact form of the form factor is available [49, 50, 51, 52]. In the thermodynamic limit ($N,L\rightarrow\infty$) the form factor decays super-exponentially in the system size as $\sim e^{-L^{2}}$ if the two sates $I,I^{\prime}$ correspond to thermodynamically distinct states, i.e. a macroscopic number of quasimomenta differ between the two [53]. A direct consequence is that the high-dimensional quantum network disconnects in most directions. Therefore, we expect that the MBDL still survive in the thermodynamic limit, at small values of $K$ and any value of $g$.